OpenStudy (anonymous):

Benzene has a normal freezing point of 5.51oC. The solution of 1.25 g of an unknown compound to 85.0 g of benzene produces a solution with a freezing point of 5.52oC. What is the molar mass of the unknown compound?

OpenStudy (aaronq):

use the formula \(\Delta = i*m*K_f\) substitute in \(m=\dfrac{n_{solute}}{kg_{solvent}}\) and \(n_{solute}=\dfrac{m_{solute}}{M_{solute}}\)

OpenStudy (aaronq):

\(\Delta T\)*****